Sachs Brands’ defined benefit pension plan specifies annual retirement benefits equal to: 1.6% × service years × final year’s salary, payable at the end of each year. Angela Davenport was hired by Sachs at the beginning of 2004 and is expected to retire at the end of 2038 after 35 years’ service. Her retirement is expected to span 18 years. Davenport’s salary is $90,000 at the end of 2018 and the company’s actuary projects her salary to be $240,000 at retirement. The actuary’s discount rate is 7%. (FV of $1, PV of $1, FVA of $1, PVA of $1, FVAD of $1 and PVAD of $1) (Use appropriate factor(s) from the tables provided.)

Required:

2. Estimate by the projected benefits approach the amount of Davenport's annual retirement payments earned as of the end of 2018.

3. What is the company's projected benefit obligation at the end of 2018 with respect to Davenport? (Do not round intermediate calculations. Round your final answer to nearest whole dollar.)

4. If no estimates are changed in the meantime, what will be the company's projected benefit obligation at the end of 2021 (three years later) with respect to Davenport? (Do not round intermediate calculations. Round your final answer to nearest whole dollar.)

Using the same question from above:

At the beginning of 2019, the pension formula was amended to:
1.75% × Service years × Final year's salary
The amendment was made retroactive to apply the increased benefits to prior service years.

Required:
1. What is the company's prior service cost at the beginning of 2019 with respect to Davenport after the amendment described above?
2. Since the amendment occurred at the beginning of 2019, amortization of the prior service cost begins in 2019. What is the prior service cost amortization that would be included in pension expense?
3. What is the service cost for 2019 with respect to Davenport?
4. What is the interest cost for 2019 with respect to Davenport?
5. Calculate pension expense for 2019 with respect to Davenport, assuming plan assets attributable to her of $150,000 and a rate of return (actual and expected) of 10%.
Answer #1

Part 1. -- The company's prior service cost is calculated as below --

Prior service cost = Change in % x service years x Final year's salary

= 0.15% (1.75-1.60) x 15 years x $240000

= $5400

Part 3. Service cost for 2019 with respect to Davenport --

$3,840 × 10.05909* = $38,627
$38,627 × .25842** = $9,982
* present value of an ordinary annuity of $1: n=18, i=7%
** present value of $1: n=20, i=7%

PART 4 -- Interest cost for 2019 with respect to Davenport?

1.6% × 14 × $240,000 = $53,760
$53,760 × 10.05909* = $540,777
$540,777 × .24151** = $130,603
= $130,603 × 7% = $9,142